# Making Differential Privacy Easier to Use for Data Controllers and Data Analysts using a Privacy Risk Indicator and an Escrow-Based Platform

### v2

@@ 3.3. Bounding Privacy Loss Across Queries
In practice, analysts often want to run multiple queries. Combining these queries’ output will reveal new information, eventually breaking privacy as per the fundamental Law of Information Recovery (Dwork et al., 2014). In today’s deployment of DP, controllers need to set a total privacy budget on the dataset a priori and stop accepting queries when the cumulative $\epsilon$ exceeds the budget (Dwork et al., 2019). The need for setting a total budget poses a critical challenge for controllers:
@@
The solution. Instead of quantifying the maximum acceptable privacy loss, which is hard, the algorithm keeps track of the already consumed budget on a dataset, i.e., the sum of $\epsilon$ of all previously answered queries; we call it $\epsilon_{c}$. For an incoming query, the algorithm truncates the range of $\epsilon$ values to consider only those larger than $\epsilon_{c}$ (line 1). The algorithm rejects the query if it fails to find a suitable $\epsilon$ larger than $\epsilon_{c}$, meaning the controller’s privacy preferences cannot be met for any possible $\epsilon$.
@@
Considering analysts’ preferences over queries. A natural tension exists between analysts and controllers: analysts want more accurate output, whereas controllers want to protect individuals’ privacy by releasing noisy output with degrading accuracy. Because it gets increasingly harder to find a suitable $\epsilon$ after answering multiple queries as $\epsilon_{c}$ keeps increasing, analysts may need to prioritize queries of which they need more accurate output. One approach to incorporate analysts’ preferences when finding $\epsilon$ is to allow them to assign weights (i.e., priority scores) to their queries, then controllers can adjust the parameter $p$ in the algorithm based on analysts’ preferences, e.g., they can set a higher $p$ for high-priority queries so the algorithm can find larger $\epsilon$ with more accurate output. Nevertheless, solving related problems, such as how to communicate analysts’ preferences to controllers, and how to help controllers decide an appropriate $p$ that takes both analysts’ preferences and their own privacy preferences into account, are outside the scope of the paper but remain interesting directions to explore.
@@ 3.4. End-to-End DP Guarantee
Implications of choosing $\epsilon$ adaptively. First of all, if the algorithm has only answered one query, releasing the query output to the analyst satisfies DP and the total privacy loss is bounded by the $\epsilon$ that is chosen by the algorithm. However, if the algorithm answered multiple queries (for the same dataset), for each query $\epsilon$ is chosen adaptively because it is chosen based on previously answered queries — a suitable $\epsilon$ needs to be larger than $\epsilon_{c}$. Then the total privacy loss is not bounded by any fixed constant since the algorithm does not require setting a total privacy budget. This brings the question of whether releasing multiple query outputs computed using adaptively chosen $\epsilon$ causes additional privacy leakage.
Privacy odometer in fully-adaptive composition. To answer this question, we construct a privacy odometer (Rogers et al., 2016) for the algorithm, which provides a probabilistic running upper bound on the realized privacy loss after answering each query. The privacy odometer is proposed to provide privacy guarantees in fully-adaptive composition (Rogers et al., 2016), in which the privacy parameters are chosen adaptively as a function of previous interactions.
Theorem 3.4 in  (Rogers et al., 2016) has proved that there is no additional leakage to choose privacy parameters adaptively if the privacy bound is computed using sequential composition. Therefore, the privacy odometer for the Find-$\epsilon$-from-PRI algorithm under sequential composition (Dwork et al., 2014) can simply be defined as:
@@
The privacy odometer gives the same privacy bound as sequential composition; since the algorithm uses $\epsilon$-DP mechanism, $\delta_{g}$ (the global failure parameter in $(\epsilon,\delta)$-differential privacy) can be set to 0. The privacy odometer updates its running upper bound on privacy loss to $\epsilon_{c}$ after the algorithm answers each query, and releasing the composed output of all previously answered queries (on the same dataset) satisfies $\epsilon_{c}$-differential privacy.
@@
It is worth noting that although  (Rogers et al., 2016) showed the privacy odometer gives a worse bound on privacy loss than the standard advanced composition theorem (Dwork et al., 2010), more recently (Whitehouse et al., 2023) proved that in the fully-adaptive setting the privacy odometer can achieve the same asymptotic bound as advanced composition under certain conditions. For simplicity, we define the privacy odometer of our algorithm for sequential composition instead of advanced composition.
3.5. Discussion and Extensions
Algorithm complexity and optimizations. The algorithm has a worst-case time complexity of $O(nm)$, where $n$ is the number of records in the dataset, and $m$ is the number of $\epsilon$ values we test. Computing the PRI of each individual requires computing the query output of each neighboring dataset (line 1 - 1). When the number of individual records is high, this becomes a performance bottleneck: using memoization saves computation if the number of unique projected records is smaller than $n$. In addition, the query computation can be parallelized since each computation is independent of the other, i.e., the problem is embarrassingly parallel. Thus, we implement multi-processing and compute the query output for a chunk of records in each process, and we show in Section 6.3.1 that the algorithm is both efficient and scalable.
Communicating Privacy-Accuracy Trade-off. In some circumstances, the controller may want to consider the accuracy of the released output when deciding on the level of privacy protection they want to offer each individual. Prior work such as DPComp (Hay et al., 2016) and ViP (Nanayakkara et al., 2022) have explored different ways to visualize the privacy-accuracy trade-off in DP. The PRI complements the existing interfaces by providing more information on the relationship between output accuracy and individual privacy implications.
4. Helping Analysts Interpret Utility
In this section, we concentrate on solutions to help analysts interpret the utility of DP output. We revisit Challenge 2 (Section 4.1) and present a solution we call utility signaling protocol (Section 4.2).
4.1. Challenge Overview and Key Insight
Controllers work on a platform that has access to the individuals’ data and that is responsible for answering analysts’ queries using DP. An analyst submits a set of queries $Q=\{q\mid q\in\mathcal{Q}\}$ ($\mathcal{Q}$ is the query space) to the platform, and (if a suitable $\epsilon$ is found for each query) receives a set of noisy query output $S=\{\mathcal{M}(x,q,\epsilon)\mid x\in\mathcal{X}^{n},q\in Q\}$ ($\mathcal{M}$ is a DP mechanism and $\mathcal{X}$ is the data universe). A well-known challenge of using DP is that analysts do not know the utility of $S$ (Hotz et al., 2022). Concretely, analysts do not know the effect of using $S$ on their downstream task, which is the purpose for analysts to query the data in the first place.
Example 1/3. An analyst queries census data at the county level. The set of query output $S$ is aggregate data on population count, education level, and housing prices that are released using DP. The analyst uses $S$ to train a binary classifier to predict the income level of each county and obtains an accuracy of 68%. The analyst does not know if the (low) accuracy is due to the effect of DP or inappropriate training data.
The difficulty in interpreting the utility of $S$ makes it challenging for analysts to understand how the DP output will impact their downstream tasks; this is one of the core issues surrounding the discussion of the use of DP for the US Census (Ruggles et al., 2019; Santos-Lozada et al., 2020; Kenny et al., 2021; ABOWD, [n.d.]).
Observation. Analysts use the DP output $S$ for some downstream task, such as a machine learning or inference task. Downstream tasks often have an associated evaluation function, $g(\cdot)\rightarrow\mathbb{R}$, which returns a number indicating the quality of the downstream task when evaluated on a certain input. In the example above, the downstream task is to train a binary classifier, the evaluation function measures the quality of the classifier trained on DP output and returns 68% as the quality indication (accuracy in this case).
Insight: signaling DP effect. Analysts would learn the impact of DP if they had access to both the result of the evaluation function on the DP output, $g(S)$ [^2] We slightly abuse notation for simplification. and the raw output, $g(S^{\prime})$, where $S^{\prime}=\{q(x)\mid x\in\mathcal{X}^{n},q\in Q\}$ (the set of query output without using DP).
Example 2/3. Besides having access to the quality of the classifier on the DP census aggregates (68%), the analyst also gains access to the quality of the classifier if it had been trained on the raw census aggregates $S^{\prime}$, say 80%. By comparing 68% ($g(S)$) and 80% ($g(S^{\prime})$), the analyst learns the effect of added DP noise on their downstream task.
Gaining access to $g(S)$ and $g(S^{\prime})$ helps analysts interpret the impact of DP on their downstream tasks. The remaining challenge is how to provide $g(S^{\prime})$ to the analysts without giving them access to $S^{\prime}$, which is the raw query output and cannot be released to the analyst without violating the privacy of individuals in the dataset.
4.2. The utility signaling protocol
Rather than the platform directly sending $S^{\prime}$ to the analyst, which would break individuals’ privacy, the analyst submits their downstream task and associated evaluation function $g(\cdot)$ to the platform. Upon successful inspection by the controller, the platform executes the downstream task and evaluation function on behalf of the analyst, producing $g(S)$ and $g(S^{\prime})$. Last, if the controller approves, the platform releases both $g(S)$ and $g(S^{\prime})$ to the analyst, thus signaling the impact of DP on the analysts’ task without revealing $S^{\prime}$.
Example 3/3. Instead of only submitting the queries to the platform, with the utility signaling protocol, the analyst submits their queries, the downstream task, and the evaluation function. The controller inspects each and allows the platform to signal the impact of using DP to the analyst by releasing both utility results 68% and 80%.
4.2.1. Protocol Step by Step
Figure 4. End-to-end example of utility signaling protocol. We illustrate the protocol in the central DP deployment scenario where the controller has access to the raw data.
Figure 4 illustrates the utility signal protocol, which we describe next:
0. Platform. The platform has access to the dataset $x$ of individuals. It also has the capacity to run queries using DP. It may optionally implement the Find-$\epsilon$-from-PRI algorithm presented in Section 3.2333While we envision the algorithm being deployed to help controllers, both contributions are orthogonal.
1. Analyst submits queries to the platform. The platform notifies the controller.
2. Controller inspects queries. The controller inspects the analyst’s queries and allows the platform to execute the queries as long as a suitable $\epsilon$ is selected. The platform withholds both the set of DP output $S$ and raw output $S^{\prime}$ without releasing anything to the analyst by default.
3. Platform releases $S$ upon controller’s approval. The platform notifies the controller after computing $S$ and $S^{\prime}$. The controller gives permission for the platform to release $S$ to the analyst.
Up to this point, the protocol provides the analyst with the same information they can access in a typical central DP deployment. Now, we extend the protocol to implement the utility signaling component to help the analyst understand the impact of DP on their downstream task.
4. Analyst submits downstream task and evaluation function $g(\cdot)$ to the platform. The platform notifies the controller.
5. Controller inspects downstream task and evaluation function. The controller inspects the implementation of the analyst’s downstream task and evaluation function $g(\cdot)$, and decides whether it is appropriate to release $g(S)$ and $g(S^{\prime})$. With the controller’s approval, the platform executes the downstream task and evaluation function $g(\cdot)$ on both $S$ and $S^{\prime}$, withholding the results by default.
6. Platform releases $g(S)$ and $g(S^{\prime})$ upon controller’s approval. The controller can inspect $g(S^{\prime})$ one last time to ensure it is appropriate to release; $g(S)$ preserves DP since it is a post-processing (Dwork et al., 2014) step of DP output $S$. If approved, the platform releases $g(S)$ and $g(S^{\prime})$ to the analyst, who learns the impact of DP on their downstream task without ever gaining access to $S^{\prime}$ directly.
4.2.2. Privacy implications of releasing $g(S^{\prime})$
Any data release leaks information. When releasing an output that is subject to DP, such as $g(S)$, controllers understand the consequences on the individual’s privacy. Conversely, when releasing output not subject to DP, such as $g(S^{\prime})$, controllers must assess whether doing so maintains individuals’ privacy. That ultimately depends on the dataset $x$, the output $S^{\prime}$, the implementation of the downstream task and evaluation function, $g(\cdot)$, and the adversary model. The scope of the utility signaling protocol is to enable analysts to understand the impact of using DP, but not on understanding, controlling, and determining when it is safe to release $g(S^{\prime})$, which is ultimately at the discretion of the controllers and an interesting avenue of future work.
5. Escrow-Based DP Platform
We curate a set of requirements for a DP platform that implements the Find-$\epsilon$-from-PRI algorithm and the utility signaling protocol in Section 5.1. In Section 5.2, we describe the implementation of a platform built on top of a data escrow that meets these requirements.
5.1. Platform Requirements
The solutions to Challenge 1 (the Find-$\epsilon$-from-PRI algorithm), to Challenge 2 (the utility signaling protocol), and the nature of processing over private data call for a platform that meets the following requirements:
  • R1. Control dataflows: Given the three relevant entities (analyst, platform, and controller), a dataflow indicates data movement between any two of those entities. A requirement for the platform is to ensure that only permitted dataflows take place. Concretely, controllers indicate what dataflows take place, and the platform enforces the controller’s decision. In the context of our work, analysts can access for example the DP output $S$ after the controller approves, and they cannot access PRIs or any other intermediates not explicitly approved by the controller.
  • R2. Secure platform: Adversarial agents may break the requirement of controlling dataflows by breaking into the platform and forcing unintended dataflows. This is an assumed risk in the central DP model, which can be reduced by using a local DP deployment (Near et al., 2021). Unfortunately, local DP reduces the accuracy of the output compared to central DP. A requirement for the platform is to ameliorate the risk of unintended dataflows.
  • R3. Protect against adversarial downstream and evaluation functions: Adversarial analysts could submit downstream tasks and evaluation functions crafted to release more information than needed. While controllers must inspect these functions before agreeing to release the output and they are the ultimate arbiter, a requirement of the platform is to facilitate the controllers’ job.
  • R4. High Performance: The Find-$\epsilon$-from-PRI algorithm increases processing overhead. Simultaneously, security safeguards to address R2 are often at odds with high-performance computation. Together, these two facts risk making the platform impractical. A requirement is to implement the solutions to Challenge 1, Challenge 2, and meet the requirements R1, R2, and R3 while maintaining low overhead.
We now explain the implementation of a platform that addresses these three requirements.
5.2. Platform Design and Implementation
To demonstrate the feasibility of the solutions to Challenge 1 and Challenge 2, we describe how to implement the solutions above on a platform that leverages a data escrow (Xia et al., 2022) architecture.
Data escrow overview. A data escrow is a system that acts as an intermediary for storing, processing, and releasing data. Different entities interact within the escrow. Controllers register data and control the dataflows. The data is never stored or computed in the clear using end-to-end encryption. Analysts run computations on registered data in the escrow. For any data or its derived data product to leave the escrow, controllers have to create a policy on the escrow. The policies are tamper-proof since the escrow validates the identity of the controller who creates the policy and the policy remains encrypted, thus ensuring the controller fully controls the dataflow.
Meeting R1. No raw data (e.g., $x$) or derived data product (e.g., PRIs, $g(S)$, $g(S^{\prime})$) leave the escrow without an explicit policy that permits that. Policies then determine what dataflows take place. Controllers are in charge of writing policies.
Meeting R2. Data is encrypted end-to-end, even during computation. Adversaries that break into the platform cannot read plain-text data. In other words, no unintended dataflows take place.
Meeting R3. The platform can let analysts submit their downstream task and associated evaluation function, and put the burden on the controller to verify the implementation of those functions. Or it can offer standard implementation of downstream tasks and evaluation functions and let analysts call these, thus requiring scrutinizing the functions once and using them many times. The data escrow permits the deployment of programs expressing a desired functionality. In specific DP scenarios, e.g., the US Census, it is conceivable that one could provide programs implementing (at least) the more standard functionalities and query interfaces, and rely on controllers to scrutinize only one-off tasks.
Meeting R4. End-to-end data encryption (necessary to meet R2) can be achieved in many ways; for instance, DP platforms leveraging Secure Multiparty Computation (Narayan and Haeberlen, 2012; Bater et al., 2016, 2018; Wang et al., 2022; Roth et al., 2019), Homomorphic Encryption (Agarwal et al., 2018; Roy Chowdhury et al., 2020), Oblivious RAM (Bogatov et al., 2021) and Trusted Execution Environment (Nguyen et al., 2020). Our evaluation prototype uses AMD Secure Encrypted Virtualization (Kaplan et al., 2016) because it permits running unmodified software while maintaining the integrity and confidentiality of memory contents. The escrow design protects access to external resources (e.g., network and disk), ensuring data is encrypted transparently to any application. As we show in the evaluation (Section 6.2), this solution virtually removes the processing overhead. In addition, we implement a series of optimizations to the Find-$\epsilon$-from-PRI algorithm to demonstrate its efficiency and scalability when run in the escrow (Section 6.3.1).
@@ 6. Evaluation
Our evaluation shows our solutions address Challenge 1 to help controllers, and Challenge 2 and 3 using an escrow-based platform. We answer the following research questions:
@@
[…]
  • Microbenchmarks: We conduct experiments to show that
    Find-$\epsilon$-from-PRI algorithm is efficient and scalable. In addition, we study the effect of varying parameter $p$ in the algorithm.
6.1. RQ1: Using PRI To Choose Epsilon
@@ 6.1.2. Study procedure
For the treatment group, the participant is presented with the same table, but with an additional column showing the range of PRIs for each $\epsilon$ (e.g., a box drawn from the minimum to maximum PRI); we also provide a brief explanation of the PRI. Note that this is the only change between the control and treatment groups to ensure we isolate the intervention. A follow-up question will be included for the treatment group that asks the participant whether they used the PRIs when choosing $\epsilon$; we include these to ensure that, if we observe an effect, we can attribute the effect to the use of PRIs.
@@ 6.1.3. Results
Figure 6 shows the $\epsilon$ chosen by participants based on their familiarity. We find similar patterns. Participants in the treatment group tend to choose either 2 or 1.5, regardless of their familiarity with DP, whereas participants in the control group chose $\epsilon$ with greater variance. This suggests participant’s a priori knowledge of DP does not transfer to making better choices of $\epsilon$; on the other hand, showing PRIs helped participants make more consistent choices whether or not they know about DP previously.
How do different ways of presenting PRIs affect participants’ decisions when choosing $\epsilon$? Since this user study relies on the participants inspecting and interpreting the PRIs with their own eyes to choose $\epsilon$ instead of using the algorithm to automatically find one, how we present the PRIs to the participants can greatly affect their choices. We explored an alternative way of presenting PRIs by using histograms. To analyze the effect of using histograms, we recruited another group of participants - 20 participants passed the quiz, and asked them to complete the same survey questions in the treatment group, with the only change being presenting histograms of PRIs instead of range plots.
@@ 6.1.4. Threats to validity
6.2. RQ2: DP On Escrow-based Platform
[…]
6.2.1. Baselines for comparison
We evaluate and compare the performance of the four following baselines:
  • Data Escrow: We implemented an escrow application that takes a SQL query as input and applies the Laplace Mechanism to release the query output. Optionally, if the parameter $\epsilon$ is not given, it will use the Find-$\epsilon$-from-PRI algorithm to select a suitable $\epsilon$. We simulate the two entities (controller and analyst) in the escrow. A policy is first created by the controller to allow the query output to be released to the analyst.
[…]
  • Standalone with DP: This baseline runs the given query and applies the Laplace Mechanism to the query output in a central platform. The dataset is in plaintext. This represents the central DP deployment scenario without the overhead of additional security measures.
  • Standalone without DP: This baseline simply runs the given query without DP in a central platform. The dataset is in plaintext. We include this baseline to show the overhead of using DP.
[…]
6.2.2. Setup
We ran the experiments on n2d-standard-8 instances with 8 virtual CPUs (AMD Rome) and 64GB memory from Google Compute Engine. We used AMD Secure Encrypted Virtualization (SEV) when running the data escrow. We used Ubuntu 20.04 and Python 3.8 for all baselines.
6.2.3. Dataset
[…]
6.2.4. Queries
We run the 3 queries independently for all baselines:
  • Q1: Number of people between age 20 and 30
  • Q2: Number of people by each race
  • Q3: Number of females who have income greater than 50K
6.2.5. Results
[…]
Figure 8. Average query time (in log scale) over 10 runs
[…]
6.3. Microbenchmarks
We evaluate the runtime and scalability of the Find-$\epsilon$-from-PRI algorithm in Section 6.3.1 and study how different values of the parameter $p$ affect the $\epsilon$ chosen in Section 6.3.2.
6.3.1. Is Find-$\epsilon$-from-PRI practical?
Since the algorithm can run multiple times if there are multiple queries , the algorithm needs to be efficient so running it multiple times will still be practical. We ran the algorithm in the data escrow to further demonstrate its performance for computation-heavy tasks.
Datasets. We controlled the experiment by generating datasets of size 1000, 10000, 100000, and 1 million based on the Adult dataset. We sampled records for datasets of sizes 1000 and 10000. For datasets of size 100000 and 1 million, we replicated and added random noise to the replicated records to ensure that they were not the same copy of each other. We included all attributes in the dataset.
Queries. We ran the algorithm to find $\epsilon$ for the five SQL queries below. We chose a diverse set of queries with attributes of different domain sizes and types. The queries are run independently.

### v3

@@ 5.2. Bounding Privacy Loss Across Queries
The preceding discussion is valid for one query. In practice, analysts often want to run multiple queries. Combining these queries’ output will reveal new information, eventually breaking privacy as per the fundamental Law of Information Recovery (Dwork et al., 2014). In today’s deployment of DP, controllers need to set a total privacy budget on the dataset a priori and stop accepting queries when the cumulative $\epsilon$ exceeds the budget (Dwork et al., 2019). The need for setting a total budget poses a critical challenge for controllers:
@@
The solution. Instead of quantifying the maximum acceptable privacy loss, which is difficult, the algorithm keeps track of the already consumed budget on a dataset, i.e., the sum of $\epsilon$ or $\epsilon+\epsilon_{svt}$ of all previously answered queries for Find-$\epsilon$-from-PRI and Find-and-release-$\epsilon$-from-PRI respectively; we call it $\epsilon_{c}$. For an incoming query, the algorithm truncates the range of $\epsilon$ values to consider only those larger than $\epsilon_{c}$ (line 1). The algorithm rejects the query if it fails to find a suitable $\epsilon$ larger than $\epsilon_{c}$, meaning the controller’s privacy preferences cannot be met for any possible $\epsilon$.
@@
Considering analysts’ preferences over queries. A natural tension exists between analysts and controllers: analysts want more accurate output, whereas controllers want to protect individuals’ privacy by releasing noisy output. Because it gets increasingly harder to find a suitable $\epsilon$ after answering multiple queries as $\epsilon_{c}$ keeps increasing, analysts may need to prioritize queries of which they need more accurate output. One approach to incorporate analysts’ preferences when finding $\epsilon$ is to allow them to assign weights (i.e., priority scores) to their queries, then controllers can adjust the parameter $p$ or $\tau$ in each algorithm based on analysts’ preferences, e.g., they can set a higher $p$ or $\tau$ for high-priority queries so the algorithm can find larger $\epsilon$ with more accurate output. Nevertheless, solving related problems, such as how to communicate analysts’ preferences to controllers, and how to help controllers decide an appropriate $p$ or $\tau$ that takes both analysts’ preferences and their own privacy preferences into account, are outside the scope of the paper but remain interesting directions to explore.
@@ 5.3. End-to-End DP Guarantee
Implications of choosing $\epsilon$ adaptively. If the algorithm has only answered one query, releasing the query output to the analyst satisfies DP and the total privacy loss is bounded by the $\epsilon$ or $\epsilon+\epsilon_{svt}$ for Find-$\epsilon$-from-PRI and Find-and-release-$\epsilon$-from-PRI respectively. However, if the algorithm answered multiple queries (for the same dataset), for each query $\epsilon$ is chosen adaptively because it is chosen based on previously answered queries — a suitable $\epsilon$ needs to be larger than $\epsilon_{c}$. Then the total privacy loss is not bounded by any fixed constant since the algorithm does not require setting a total privacy budget. This brings the question of whether releasing multiple query outputs computed using adaptively chosen $\epsilon$ causes additional privacy leakage.
Privacy odometer in fully-adaptive composition. To answer this question, we construct a privacy odometer (Rogers et al., 2016), which provides a probabilistic running upper bound on the realized privacy loss after answering each query. The privacy odometer is proposed to provide privacy guarantees in fully-adaptive composition (Rogers et al., 2016), in which the privacy parameters are chosen adaptively as a function of previous interactions.
In short, there is no additional leakage to choose privacy parameters adaptively if the privacy bound is computed using sequential composition. Therefore, the privacy odometer for Find-$\epsilon$-from-PRI under sequential composition (Dwork et al., 2014) is defined as:
@@
The privacy odometer for Find-and-release-$\epsilon$-from-PRI can be defined using the same principle, which we omit. Essentially, the privacy odometer updates its running upper bound on privacy loss to $\epsilon_{c}=\sum_{i=1}^{i=k}\epsilon_{i}$ after the algorithm answers the $k$-th query, and releasing the composed output of all previously answered queries (on the same dataset) satisfies $(\epsilon_{c},\delta_{g})$-DP (we can set $\delta_{g}=\sum_{i=1}^{i=k}\delta$).
@@
It is also worth noting that under certain conditions (Whitehouse et al., 2023), the privacy odometer achieves the same asymptotic bound as advanced composition (Dwork et al., 2010). For simplicity, we define the privacy odometer of our algorithm for sequential composition instead of advanced composition, and we leave further improvements for future work.
@@ 6. Evaluation
We answer the following research questions:
@@
  • RQ2: Are our proposed algorithms, Find-$\epsilon$-from-PRI and
    Find-and-release-$\epsilon$-from-PRI efficient and scalable?
  • Microbenchmarks: We study the effect of varying parameter $p$ in Find-$\epsilon$-from-PRI, and parameter $\tau$ in Find-and-release-$\epsilon$-
    from-PRI
    , and we compare the chosen $\epsilon$ values by both algorithms.
6.1. RQ1: Does PRI Help Controllers Choose $\epsilon$?
@@ 6.1.2. Study procedure
For the treatment group, the participant is presented with the same table, but with an additional column showing the range of PRIs for each $\epsilon$ (e.g., a box drawn from the minimum to maximum PRI); we also provide a brief explanation of the PRI. Note that this is the only change between the control and treatment groups to ensure we isolate the intervention. A follow-up question will be included for the treatment group that asks the participant whether they used the PRIs when choosing $\epsilon$; we include it to ensure that, if we observe an effect, we can attribute the effect to the use of PRIs.
@@ 6.1.3. Results
Figure 6 shows the $\epsilon$ chosen by participants based on their familiarity. We find similar patterns. Participants in the treatment group tend to choose either 2 or 1.5, regardless of their familiarity with DP, whereas participants in the control group chose $\epsilon$ with greater variance. This suggests participant’s a priori knowledge of DP does not transfer to making better choices of $\epsilon$; on the other hand, showing PRIs helped participants make more consistent choices whether or not they knew about DP previously.
How do different ways of presenting PRIs affect participants’ decisions when choosing $\epsilon$? Since this user study relies on the participants inspecting and interpreting the PRIs with their own eyes to choose $\epsilon$ instead of using the algorithm to automatically find one, how we present the PRIs to the participants can greatly affect their choices. We explored an alternative way of presenting PRIs by using histograms. To analyze the effect of using histograms, we recruited another group of participants - 20 participants passed the quiz, and we asked them to complete the same survey questions as in the treatment group except that the survey presents histograms of PRIs instead of range plots.
@@ 6.1.4. Threats to validity
6.2. RQ2: Are Both Algorithms Practical?
Since both Find-$\epsilon$-from-PRI and Find-and-release-$\epsilon$-from-PRI algorithms can run multiple times if there are multiple queries, the algorithms need to be efficient so running it multiple times will still be practical. Thus, we evaluate both algorithms’ runtime for different queries and dataset sizes.
Datasets. We controlled the experiment by generating datasets with 1000, 10000, 100000, and 1 million records based on the Adult dataset (uci, [n. d.]). We sampled records for datasets of sizes (number of records) 1000 and 10000. For datasets of size 100000 and 1 million, we replicated and added random noise to the replicated records to ensure that they were not the same copy of each other. We included all attributes in the dataset.
Queries. We ran both algorithms for the five SQL queries below. We chose a diverse set of queries with attributes of different domain sizes and types. The queries are run independently.